OpenStudy (anonymous):

What is the solution set of |x – 6| = 8? 14 14 or –14 14 or –2 14 or 2

OpenStudy (anonymous):

do same two equations where on one side the eight remains positive and other equation the 8 is a negative.

OpenStudy (anonymous):

you remove the "l l" symbols THEY DON'T MATTER.

OpenStudy (anonymous):

|x – 6| = 8 x – 6 = ±8 x = 6 ± 8 = 6+8 or 6-8 = 14 or -2